Circular No. 8376

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004di IN UGC 10097 Independent discovery reports of an apparent supernova have been re-

ceived from M. Armstrong (Rolvenden, U.K., 0.35-m reflector) and R. Arbour (South Wonston, U.K., 0.3-m reflector). Armstrong reports the new object’s position as α = 15^{h}55^{m}41^{s}.02, δ = +47^{o}51′49′′.1 (equinox 2000.0), which is ≈ 23′′ west and 11′′ south of the center of UGC 10097; Arbour’s measured position end figures are 40^{s}.98, 48′′.8. Unfiltered CCD magnitudes of SN 2004di: July 14 UT, [19.0 (Armstrong); 18.922, [17.5 (Arbour); 26.903, 17.0 (Armstrong); 27.891, 17.0 (Armstrong); 27.911, 16.8 (Arbour). Nothing was visible at this position on Palomar Sky Survey images from 1989 and 1993 (red limiting mag 20.8; blue limiting mag 22.5).

SUPERNOVA 2004dg N. Elias de la Rosa, S. Benetti, H. Navasardyan, F. Bufano, A. Haru-

tyunyan, G. Blanc, and M. Turatto, Osservatorio Astronomico di Padova; A. Pastorello, Osservatorio Astrofisico di Arcetri; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, report that inspection of a spectrogram, obtained on July 21.87 UT with the Asiago 1.8-m telescope (+ AFOSC; range 355–780 nm, resolution 2.4 nm), indicates that SN 2004dg (cf. IAUC 8375) is a type-II supernova, ∼ 1 month after explosion. The spectrum is dominated by P-Cyg lines of H I, Ca II H and K, Fe II, and Ti II. A narrow Hα feature (from a nearby H II region) is visible on the top of the broad emission, from which the observers deduce a recession velocity of ∼ 1720 km/s, which in turn is substantially higher than the galaxy’s recession velocity given by NED (1359 km/s). The expansion velocities deduced from the Hα and Hβ absorptions are ∼ 8270 and 6760 km/s, respectively.

SUPERNOVAE 2004ct AND 2004dh T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 340–730 nm) of SN 2004ct (cf. IAUC 8361), obtained by M. Calkins on July 22.28 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova several weeks past maximum. A spectrogram of SN 2004dh (cf. IAUC 8375), obtained by Calkins on July 23.44, shows it to be a type-II supernova. The spectrum consists of a blue continuum with P-Cyg lines of hydrogen. Adopting the NED recession velocity of 5794 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 9500 km/s.

2004 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green